# Right on Time: Revising Time Series Models by Constraining their Explanations

### v3

@@ 1 Introduction
Time series data is ubiquitous in our world today. Everything that is measured over time generates some form of time series, for example, energy load (Koprinska et al., 2018), sensor measurements in industrial machinery (Mehdiyev et al., 2017) or recordings of traffic data (Ma et al., 2022). Various neural models are often applied to complex time series data (Ruiz et al., 2021; Benidis et al., 2023). As in other domains, these can be subject to confounding factors ranging from simple noise or artifacts to complex shortcut confounders (Lapuschkin et al., 2019). Intuitively, a confounder, also called “Clever-Hans” moment, can be a pattern in the data which is not relevant for the task, but correlates with it during model training. A model can incorrectly pick up on this confounder (cf. Fig. 1) and use it instead of the relevant features to e.g. make a classification. A confounded model does not generalize well to data without the confounder, which is a problem when employing models in practice Geirhos et al. (2020). For time series, confounders and their mitigation have yet to receive attention, where existing works make specific assumptions about settings and data (Bica et al., 2020).
@@
In particular, the mitigation of shortcut confounders, i.e., spurious patterns in the training data used for the prediction, is essential. If a model utilizes confounding factors in the training set, its decision relies on wrong reasons, which makes it fail to generalize to data without the confounder. Model explanations play a crucial role in uncovering confounding factors, but they are not enough on their own to address them. While an explanation can reveal that the model relies on incorrect factors, it does not alter the model’s outcome. To change this, we introduce Right on Time (RioT), a new method following the core ideas of explanatory interactive learning (XIL) (Teso and Kersting, 2019), i.e., using feedback on explanations to mitigate confounders. RioT uses traditional explanation methods like Integrated Gradients (IG) (Sundararajan et al., 2017) to detect whether models focus on the right or the wrong time steps and utilizes feedback on the latter to revise the model (Fig. 1, left).
@@
Confounding factors in time series data are not limited to the time domain. A steady noise frequency in an audio signal can also be a confounder but cannot be pinned to a specific time step. To handle these kinds of confounders, RioT can also incorporate feedback in the frequency domain (Fig. 1, right). To further emphasize the importance of mitigating confounders in time series data, we introduce a new real-world, confounded dataset called Production Press Sensor Data (P2S). The dataset consists of sensor measurements from an industrial high-speed press, which is part of many important manufacturing processes in the sheet metal working industry. The sensor data used to detect faulty production is naturally confounded and thus causes incorrect predictions after training. Next to its immediate industrial relevance, P2S is the first time series dataset that contains explicitly annotated confounders, enabling evaluation and comparison of confounder mitigation strategies on real data.
Figure 1: Left: Expert annotations marking the wrong reasons (red) are given based on model explanations to mitigate spatial confounders. The application of RioTsp revises the model, whose adjusted explanations reveal that its output is based on the right reasons (blue) instead. Right: As confounders can also occur in the frequency domain, RioTfreq can use explanations and feedback in this domain as well. Similar to the spatial domain, it guides the model’s focus, mitigating the influence of the confounders on the target prediction.
@@ 2 Related Work
Explanations for Time Series. Within the field of explainable artificial intelligence (XAI), various techniques to explain machine learning models and their outcomes have been proposed. While many techniques originate from image or text data, they were quickly adapted to time series Rojat et al. (2021). While backpropagation- and perturbation-based attribution methods provide explanations directly in the input space, other techniques like symbol aggregations (Lin et al., 2003) or shapelets (Ye and Keogh, 2011) aim to provide higher-level explanations. For a more in-depth discussion of explanations for time series, we refer to surveys by Rojat et al. (2021) or Schlegel et al. (2019). While it is essential to have explanation methods to detect confounding factors, they alone are insufficient to revise a model. Thus, explanations are the starting point of our method, as they enable users to detect confounders and provide feedback to overcome confounders in the model. In particular, we build upon Integrated Gradients (IG) (Sundararajan et al., 2017). This method computes attributions for the input by utilizing model gradients. We selected it because of its several desirable properties like completeness or implementation invariance and its wide use, also for time series data (Mercier et al., 2022; Veerappa et al., 2022).
@@
Explanatory Interactive Learning (XIL). While not prevalent for time series data, there has been some work on confounders and how to overcome them for other domains, primarily the image domain. Most notable, there is explanatory interactive learning, which describes the general process of revising a model’s decision process based on human feedback (Teso and Kersting, 2019; Schramowski et al., 2020). Within XIL, the model’s explanations are used to incorporate the feedback back to the model, thus revising its mistakes (Friedrich et al., 2023a). XIL can be used on models that show Clever-Hans-like behavior (being affected by shortcuts in the data) to prevent them from using these shortcuts Stammer et al. (2020). Several methods apply the idea of XIL to image data. For example, Right for the Right Reasons (RRR) (Ross et al., 2017) and Right for Better Reasons (RBR) (Shao et al., 2021) use human feedback as a penalty mask on model explanations. Instead of penalizing wrong reasons, HINT (Selvaraju et al., 2019) rewards the model for focusing on the correct part of the input. Furthermore, Friedrich et al. (2023b) investigate the use of multiple explanation methods simultaneously. Although various XIL methods are often employed to address confounders in image data, their application to time series data remains unexplored. To bridge this gap, we introduce RioT, a method that incorporates the core principles of XIL to the unique characteristics of time series data.
@@
Unconfounding Time Series. Next to approaches from interactive learning, there is also some other work on unconfounding time series models. This line of work is generally based on causal analysis of the time series model and data (Flanders et al., 2011). Methods like Time Series Deconfounder (Bica et al., 2020), SqeDec (Hatt and Feuerriegel, 2024) or LipCDE (Cao et al., 2023), perform estimations on the data while mitigating the effect of confounders in covariates of the target variable. They generally mitigate the effect of the confounders through casual analysis and specific assumptions about the data generation. On the other hand, in this work we tackle confounders within the target variate, and have no further assumption besides that the confounder is visible in the explanations of the model, where these previous methods cannot easily be applied.
@@ 3 Right on Time (RioT)
Figure 2: This figure depicts the flow of explanation and revision, beginning with input data $\bm{x}$, through the model $f(\bm{x})$ to explanations $e(\bm{x})$, annotated feedback $a(\bm{x})$, and finally back to the model. IG provides the spatial model explanation, and is transformed with the FFT to obtain the frequency explanation. Expert user annotations can be applied in either or both domains. They are utilized by the right-reason loss ($\mathcal{L}_{RR}^{sp}$ and $\mathcal{L}_{RR}^{freq}$) to guide the model away from confounders in both the time and frequency domains.
The core intuition of Right on Time (RioT) is to utilize human feedback to steer a model away from wrong reasons. It follows the general structure of XIL, which has four main steps Friedrich et al. (2023a). In Select, instances for feedback and following model revision are selected. Following previous XIL methods, we select all samples by default while not necessarily requiring feedback for all of them. Afterwards, Explain covers how model explanations are generated, before in Obtain, a human provides feedback on the selected instances. Lastly, in Revise, the feedback is integrated into the model to overcome the confounders. We introduce RioT along these steps in the following (the entire process is illustrated in Fig. 2). But let us first establish some notation for the remainder of this paper.
Given is a dataset $(\mathcal{X},\mathcal{Y})$ and a model $f(\cdot)$ for time series classification or forecasting. The dataset consists of $D$ many pairs of $\bm{x}$ and $\bm{y}$. Thereby, $\bm{x}\in\mathcal{X}$ is a time series of length $T$, i.e., $\bm{x}\in\mathbb{R}^{T}$. For $K$ class classification, the ground-truth output is $\bm{y}\in\{1,\dots,K\}$ and for forecasting, the ground-truth output is the forecasting window $\bm{y}\in\mathbb{R}^{W}$ of length $W$. The ground-truth output of the full dataset is then described as $\mathcal{Y}$ in both cases. For a datapoint $\bm{x}$, the model generates the output $\hat{\bm{y}}=f(\bm{x})$, where the dimensions of $\hat{\bm{y}}$ are the same as of $\bm{y}$ for both tasks.
@@ 3.1 Explain
Given a pair of input $\bm{x}$ and model output $\hat{\bm{y}}$ for time series classification, the explainer generates an explanation $e_{f}(\bm{x})\in\mathbb{R}^{T}$ in the form of attributions to explain $\hat{\bm{y}}$ w.r.t. $\bm{x}$. For an element of the input, a large attribution value means a large influence on the output. In the remainder of the paper, explanations refer to the model $f$, but we drop $f$ from the notation to declutter it, resulting in $e(\bm{x})$. We use IG (Sundararajan et al., 2017) as an explainer, an established gradient-based attribution method. However, we make some adjustments to the base method to make it more suitable for time series and model revision (Eq. 1, further details in SubSec. A.2). In the following, we introduce the modifications to use attributions for forecasting and to obtain explanations in the frequency domain.
@@
Attributions for Forecasting. In a classification setting, attributions are generated by propagating gradients back from the model output (of its highest activated class) to the model inputs. However, there is often no single model output in time series forecasting. Instead, the model generates one output for each timestep of the forecasting window simultaneously. Naively, one could use these $W$ outputs and generate as many explanations $e^{\prime}_{1}(\bm{x}),\dots e^{\prime}_{W}(\bm{x})$. This number of explanations would, however, make it even harder for humans to interpret the results, as the size of the explanation increases with $W$(Miller, 2019). Therefore, we propose to aggregate the individual explanations by averaging in Eq. 2. Averaging attributions over the forecasting window provides a simple yet robust aggregation of the explanations. Other means of combining them, potentially even weighted based on distance of the forecast in the future are also imaginable. Overall, this allows attributions for time series classification and forecasting to be generated similarly.
@@ 3.2 Obtain
The next step of RioT is to obtain user feedback on confounding factors. For an input $\bm{x}$, a user can mark parts that are confounded, resulting in a feedback mask $a(\bm{x})\in\{0,1\}^{T}$. In this binary mask, a $1$ signals a potential confounder at this time step. Thereby, it is not necessary to have feedback for every sample of the dataset, as a mask $a(\bm{x})=(0,\dots,0)^{T}$ corresponds to no feedback. Feedback can also be given on the frequency explanation in a similar manner, marking which elements in the frequency domain are potential confounders. The resulting feedback mask $\hat{a}(\bm{x})=(\hat{a}(\bm{x})_{re},\hat{a}(\bm{x})_{im})$ can be different for the real $\hat{a}(\bm{x})_{re}\in\{0,1\}^{T}$ and imaginary part $\hat{a}(\bm{x})_{im}\in\{0,1\}^{T}$. For the whole dataset, we then have spatial annotations $A$ and frequency annotations $\hat{A}$. As the annotated feedback masks have to come from human experts, obtaining them can become costly. In many cases, however, confounders occur systematically and it is therefore possible to apply the same annotation mask to multiple samples. This can drastically reduce the number of annotations required in practice.
@@ 4.1 Experimental Setup
Data. We perform experiments on various datasets. For classification, we focus mainly on the UCR/UEA repository (Dau et al., 2018), which holds a wide variety of datasets for this task. The data originates from different domains, e.g., health records, industrial sensor data, and audio signals. We select all available datasets of a minimal size (cf. SubSec. A.3), which results in Fault Detection A, Ford A, Ford B, and Sleep. We omit experiments on the very small datasets of UCR, as these are generally less suited for deep learning (Ismail Fawaz et al., 2020). We use the splits provided by the UCR archive. For time series forecasting, we evaluate on three popular datasets of the Darts repository (Herzen et al., 2022): ETTM1, Energy, and Weather with 70%/30% train/test splits. These datasets are sufficiently large, allowing us to investigate the impact of confounding behavior in isolation without the risk of overfitting. We standardize all datasets as suggested by Wu et al. (2021), i.e., rescaling the distribution of values to zero mean and a standard deviation of one.
@@
Table 1: Applying RioT mitigates confounders in time series classification. Performance before and after applying RioT for spatial (SP Conf) and frequency (Freq Conf) confounders separately. Unconfounded represents the ideal scenario where the model is not affected by any confounder.
@@
SP Conf 1.00 ±0.00 0.74 ±0.06 1.00 ±0.00 0.71 ±0.08 1.00 ±0.00 0.63 ±0.03 1.00 ±0.00 0.10 ±0.03
+ RioTsp 0.98 ±0.01 $\bullet$ 0.93 ±0.03 0.99 ±0.01 $\bullet$ 0.84 ±0.02 0.99 ±0.00 $\bullet$ 0.68 ±0.02 0.60 ±0.06 $\bullet$ 0.54 ±0.05
Freq Conf 0.98 ±0.01 0.87 ±0.03 0.98 ±0.00 0.73 ±0.01 0.99 ±0.01 0.60 ±0.01 0.98 ±0.00 0.27 ±0.02
+ RioTfreq 0.94 ±0.00 $\bullet$ 0.90 ±0.03 0.83 ±0.02 $\bullet$ 0.83 ±0.02 0.94 ±0.00 $\bullet$ 0.65 ±0.01 0.67 ±0.05 $\bullet$ 0.45 ±0.07
@@
SP Conf 1.00 ±0.00 0.53 ±0.02 1.00 ±0.00 0.50 ±0.00 1.00 ±0.00 0.52 ±0.01 1.00 ±0.00 0.21 ±0.05
+ RioTsp 0.96 ±0.08 $\bullet$ 0.98 ±0.01 0.92 ±0.03 $\bullet$ 0.85 ±0.02 0.94 ±0.01 $\bullet$ 0.65 ±0.04 0.52 ±0.22 $\bullet$ 0.58 ±0.05
Freq Conf 1.00 ±0.00 0.72 ±0.02 1.00 ±0.00 0.65 ±0.01 1.00 ±0.00 0.56 ±0.02 0.99 ±0.00 0.24 ±0.03
+ RioTfreq 0.96 ±0.02 $\bullet$ 0.98 ±0.02 0.78 ±0.04 $\bullet$ 0.85 ±0.04 1.00 ±0.00 $\bullet$ 0.64 ±0.03 0.50 ±0.16 $\bullet$ 0.49 ±0.04

### v4

@@ 1 Introduction
Time series data is ubiquitous in our world today. Everything that is measured over time generates some form of time series, for example, energy load (Koprinska et al., 2018), sensor measurements in industrial machinery (Mehdiyev et al., 2017) or recordings of traffic data (Ma et al., 2022). Various neural models are often applied to complex time series data (Ruiz et al., 2021; Benidis et al., 2023). As in other domains, these can be subject to confounding factors ranging from simple noise or artifacts to complex shortcut confounders (Lapuschkin et al., 2019). Intuitively, a confounder, also called “Clever-Hans” moment, can be a pattern in the data that is not relevant to the task but correlates with it during model training. A model can incorrectly pick up on this confounder and use it instead of the relevant features to e.g. make a classification. A confounded model does not generalize well to data without the confounder, which is a problem when employing models in practice (Geirhos et al., 2020). For time series, confounders and their mitigation have yet to receive attention, where existing works make specific assumptions about settings and data (Bica et al., 2020).
@@
In particular, it is essential to mitigate shortcut confounders, i.e., spurious patterns in the training data used for the prediction. If a model utilizes confounding factors in the training set, its decision relies on wrong reasons and fails to generalize to unconfounded data. Model explanations play a crucial role in uncovering confounding factors, but they are not enough on their own to address them (cf. Fig. 1 I). While an explanation can reveal that the model relies on incorrect factors, it does not alter the model’s outcome. To change this, we introduce Right on Time (RioT), a new method following the core ideas of the explanatory interactive learning (XIL) paradigm (Teso & Kersting, 2019), i.e., using feedback on explanations to mitigate confounders (cf. Fig. 1 II). RioT uses traditional explanation methods like Integrated Gradients (IG) (Sundararajan et al., 2017) to detect whether models focus on the right or the wrong time steps and utilizes feedback on the latter to revise the model.
@@
However, confounding factors in time series data extend beyond the time domain. For example, a consistent noise frequency in an audio signal can act as a confounder without being tied to a specific point in time. RioT can handle these types of confounders by incorporating feedback in the frequency domain. To highlight the importance of mitigating confounders in time series data, we introduce a new real-world, confounded dataset called Production Press Sensor Data (P2S). The dataset includes sensor measurements from an industrial high-speed press, essential to many manufacturing processes in the sheet metal working industry. The sensor data for detecting faulty production is naturally confounded and thus causes incorrect predictions after training. Next to its immediate industrial relevance, P2S is the first time series dataset that contains explicitly annotated confounders, enabling the evaluation and comparison of confounder mitigation strategies on real data.
Figure 1: I: Explanations can reveal whether models rely on confounding factors in the input instead of relevant features. With RioT, a human can provide feedback on the spatial and frequency domain explanations for wrong reasons. This feedback is used to revise the model to not consider those regions. II: After revising via RioT, the model focuses on the right reasons instead.
@@ 2 Related Work
Explanations for Time Series. Within the field of explainable artificial intelligence (XAI), various techniques to explain machine learning models and their outcomes have been proposed. While many techniques originate from image or text data, they were quickly adapted to time series Rojat et al. (2021). While backpropagation- and perturbation-based attribution methods provide explanations directly in the input space, other techniques like symbol aggregations (Lin et al., 2003) or shapelets (Ye & Keogh, 2011) aim to provide higher-level explanations. For a more in-depth discussion of explanations for time series, we refer to surveys by Rojat et al. (2021) or Schlegel et al. (2019). While it is essential to have explanation methods to detect confounding factors, they alone are insufficient to revise a model. Thus, explanations are the starting point of our method, as they enable users to detect confounders and provide feedback to overcome confounders in the model. In particular, we build upon Integrated Gradients (IG) (Sundararajan et al., 2017). This method computes attributions for the input by utilizing model gradients. We selected it because of its several desirable properties like completeness or implementation invariance and its wide use, also for time series data (Mercier et al., 2022; Veerappa et al., 2022).
@@
Explanatory Interactive Learning (XIL). There is a growing field of research investigating confounding factors and how to overcome them. However, these mainly focus on visual data domains. One paradigm to overcome these confounders is using explanatory interactive learning, which describes the general process of revising a model’s decision process based on human feedback(Teso & Kersting, 2019; Schramowski et al., 2020). A key element of XIL is using the model’s explanations to incorporate human feedback, thus revising the model’s mistakes (Friedrich et al., 2023a). This is primarily done to revise models that show Clever-Hans-like behavior (being affected by shortcuts in the data) to prevent them from using these shortcuts Stammer et al. (2020). Several methods apply the XIL paradigm to image data. For example, Right for the Right Reasons (RRR) (Ross et al., 2017) and Right for Better Reasons (RBR) (Shao et al., 2021) use human feedback as a penalty mask on model explanations. Instead of penalizing wrong reasons, HINT (Selvaraju et al., 2019) rewards the model for focusing on the correct part of the input. Furthermore, Friedrich et al. (2023b) investigate the use of multiple explanation methods simultaneously. Although these methods are often employed to address confounders in image data, their application to time series data remains unexplored. To bridge this gap, we introduce RioT, a method that incorporates the core principles of XIL to the unique characteristics of time series data.
@@
Unconfounding Time Series. Next to approaches from interactive learning, there is also some other work on unconfounding time series models. This line of work is generally based on causal analysis of the time series model and data (Flanders et al., 2011). Methods like Time Series Deconfounder (Bica et al., 2020), SqeDec (Hatt & Feuerriegel, 2024) or LipCDE (Cao et al., 2023), perform estimations on the data while mitigating the effect of confounders in covariates of the target variable. They generally mitigate the effect of the confounders through causal analysis and specific assumptions about the data generation. On the other hand, in this work we tackle confounders within the target variate, and have no further assumption besides that the confounder is visible in the explanations of the model, where these previous methods cannot easily be applied.
@@ 3 Right on Time (RioT)
Figure 2: The explanation-based revision begins with input data $\bm{x}$, through the model $f(\bm{x})$ to explanations $e(\bm{x})$, annotated feedback $a(\bm{x})$, and goes finally back to the model. IG provides the spatial explanations, and transformed via FFT the frequency explanation. Annotations can be applied in either or both domains. They are utilized by the right-reason loss ($\mathcal{L}_{RR}^{sp}$ and $\mathcal{L}_{RR}^{freq}$) to guide the model away from confounders in both the time and frequency domains.
The core intuition of Right on Time (RioT) is leveraging human feedback to steer a model away from wrong reasons. The method follows the general paradigm of XIL. To make potential combinations with other methods within XIL straightforward, we develop RioT along the steps found by Friedrich et al. (2023a): Select, Explain, Obtain and Revise. In Select, samples for feedback and model revision are selected. Following previous methods, we select all samples by default but also explore using subsets of the data. Afterwards, Explain covers model explanations, before in human feedback is provided in Obtain. Lastly, in Revise, the feedback is integrated into the model to overcome the confounders. We introduce RioT along these steps in the following (as illustrated in Fig. 2). But let us first establish some notation for the remainder of this paper.
Given a dataset $(\mathcal{X},\mathcal{Y})$ and a model $f(\cdot)$ for time series classification or forecasting. The dataset consists of $D$ many pairs of $\bm{x}$ and $\bm{y}$. Thereby, $\bm{x}\in\mathcal{X}$ is a time series of length $T$, i.e., $\bm{x}\in\mathbb{R}^{T}$. For $K$ class classification, the ground-truth output is $\bm{y}\in\{1,\dots,K\}$ and for forecasting, the ground-truth output is the forecasting window $\bm{y}\in\mathbb{R}^{W}$ of length $W$. The ground-truth output of the full dataset is then described as $\mathcal{Y}$ in both cases. For a datapoint $\bm{x}$, the model generates the output $\hat{\bm{y}}=f(\bm{x})$, where the dimensions of $\hat{\bm{y}}$ are the same as of $\bm{y}$ for both tasks.
@@ 3.1 Explain
Given a pair of input $\bm{x}$ and model output $\hat{\bm{y}}$ for time series classification, the explainer generates an explanation $e_{f}(\bm{x})\in\mathbb{R}^{T}$ in the form of attributions to explain $\hat{\bm{y}}$ w.r.t. $\bm{x}$. For an element of the input, a large attribution value means a large influence on the output. In the remainder of the paper, explanations refer to the model $f$, but we drop $f$ from the notation to declutter it, resulting in $e(\bm{x})$. We use IG (Sundararajan et al., 2017) (Eq. 1) as an explainer, an established gradient-based attribution method. This method integrates the gradient along the path (using the integration variable $\alpha$) from a baseline $\bar{\bm{x}}$ to the input $x$ and multiplies the result with the difference between baseline and input. However, we make some adjustments to the base method to make it more suitable for time series and model revision, namely taking the absolute value of the difference between $\bm{x}$ and $\bar{\bm{x}}$ ( further details in SubSec. A.1). In the following, we introduce the modifications to use attributions for forecasting and to obtain explanations in the frequency domain.
@@
Attributions for Forecasting. In a classification setting, attributions are generated by propagating gradients back from the model output (of its highest activated class) to the model inputs. However, there is often no single model output in time series forecasting. Instead, the model generates one output for each timestep of the forecasting window simultaneously. Naively, one could use these $W$ outputs and generate as many explanations $e^{\prime}_{1}(\bm{x}),\dots e^{\prime}_{W}(\bm{x})$, where each $e^{\prime}_{i}(\bm{x})$ is the IG explanation using the i-th time step from the forecasting window as target instead of a classification label. This number of explanations would, however, make it even harder for humans to interpret the results, as the size of the explanation increases with $W$(Miller, 2019). Therefore, we propose to aggregate the individual explanations by averaging in Eq. 2. Averaging attributions over the forecasting window provides a simple yet robust aggregation of the explanations. Other means of combining them, potentially even weighted based on distance of the forecast in the future are also imaginable. Overall, this allows attributions for time series classification and forecasting to be generated similarly.
@@ 3.2 Obtain
The next step of RioT is to obtain user feedback on confounding factors. For an input $\bm{x}$, a user can mark parts that are confounded, resulting in a feedback mask $a(\bm{x})\in\{0,1\}^{T}$. In this binary mask, a $1$ signals a potential confounder at this time step. Thereby, it is not necessary to have feedback for every sample of the dataset, as a mask $a(\bm{x})=(0,\dots,0)^{T}$ corresponds to no feedback. Feedback can also be given on the frequency explanation in a similar manner, marking which elements in the frequency domain are potential confounders. The resulting feedback mask $\hat{a}(\bm{x})=(\hat{a}(\bm{x})_{re},\hat{a}(\bm{x})_{im})$ can be different for the real $\hat{a}(\bm{x})_{re}\in\{0,1\}^{T}$ and imaginary part $\hat{a}(\bm{x})_{im}\in\{0,1\}^{T}$. For the whole dataset, we then have spatial annotations $A$ and frequency annotations $\hat{A}$.
As the annotated feedback masks come from human experts, obtaining them can become costly. However, confounders often occur systematically, and it is thus possible to apply the same annotation mask to many samples. This can drastically reduce the number of annotations required in practice.
@@ 4.1 Experimental Setup
Data. We perform experiments on various datasets. For classification, we focus mainly on the UCR/UEA repository (Dau et al., 2018), which holds a wide variety of datasets for this task. The data originates from different domains, e.g., health records, industrial sensor data, and audio signals. We select all available datasets of a minimal size (cf. SubSec. A.2), which results in Fault Detection A, Ford A, Ford B, and Sleep. We omit experiments on the very small datasets of UCR, as these are generally less suited for deep learning (Ismail Fawaz et al., 2020). We use the splits provided by the UCR archive. For time series forecasting, we evaluate on three popular datasets from the Darts repository (Herzen et al., 2022): ETTM1, Energy, and Weather. We split the data into training and test sets using a 70/30 ratio. The training set is further divided into training and validation subsets in an 80/20 ratio, resulting in an overall train/validation/test split of 56%/14%/30%. These datasets are sufficiently large, allowing us to investigate the impact of confounding behavior in isolation without the risk of overfitting. We standardize all datasets as suggested by Wu et al. (2021), i.e., rescaling the distribution of values to zero mean and a standard deviation of one.
@@
Table 1: Applying RioT mitigates confounders in time series classification. Performance before and after applying RioT for spatial (Basesp) and frequency (Basefreq) confounders. High training and low test accuracies indicate overfitting to the confounder, which RioT successfully mitigates. Unconfounded represents the ideal scenario where the model is not affected by any confounder.
@@
Basesp 1.00 ±0.00 0.74 ±0.06 1.00 ±0.00 0.71 ±0.08 1.00 ±0.00 0.63 ±0.03 1.00 ±0.00 0.10 ±0.03
+ RioTsp 0.98 ±0.01 0.93 ±0.03 0.99 ±0.01 0.84 ±0.02 0.99 ±0.00 0.68 ±0.02 0.60 ±0.06 0.54 ±0.05
Basefreq 0.98 ±0.01 0.87 ±0.03 0.98 ±0.00 0.73 ±0.01 0.99 ±0.01 0.60 ±0.01 0.98 ±0.00 0.27 ±0.02
+ RioTfreq 0.94 ±0.00 0.90 ±0.03 0.83 ±0.02 0.83 ±0.02 0.94 ±0.00 0.65 ±0.01 0.67 ±0.05 0.45 ±0.07
@@
Basesp 1.00 ±0.00 0.53 ±0.02 1.00 ±0.00 0.50 ±0.00 1.00 ±0.00 0.52 ±0.01 1.00 ±0.00 0.21 ±0.05
+ RioTsp 0.96 ±0.08 0.98 ±0.01 0.92 ±0.03 0.85 ±0.02 0.94 ±0.01 0.65 ±0.04 0.52 ±0.22 0.58 ±0.05
Basefreq 1.00 ±0.00 0.72 ±0.02 1.00 ±0.00 0.65 ±0.01 1.00 ±0.00 0.56 ±0.02 0.99 ±0.00 0.24 ±0.03
+ RioTfreq 0.96 ±0.02 0.98 ±0.02 0.78 ±0.04 0.85 ±0.04 1.00 ±0.00 0.64 ±0.03 0.50 ±0.16 0.49 ±0.04